CURRICULUM VITAE

Name  — Krzysztof Galicki
Born  — August 10, 1958, Poland
Citizenship  — United States
Current address  — Department of Mathematics & Statistics

The University of New Mexico Albuquerque, NM 87131, U.S.A. home tel. # (505) 271-9883 office tel. # (505) 277-2024 office fax. # (505) 277-5505 e-mail: galicki@math.unm.edu

• Education

August 1987  — Doctor of Philosophy in Mathematical Physics,

SUNY@Stony Brook, Stony Brook, NY, U.S.A.

May 1987  — Master of Arts,

SUNY@Stony Brook, Stony Brook, NY, U.S.A.

April 1982  — Master’s Degree in Theoretical Physics,

University of Wrocław, Wrocław, Poland

• Area of Research  — Differential Geometry and Mathematical Physics

• Current Position

08/1999 - present  — Professor

— Department of Mathematics & Statistics,

— University of New Mexico, Albuquerque, NM, U.S.A.

• Previous Positions

08/1993 - 08/1999  — Associate Professor, Department of Mathematics

& Statistics, University of New Mexico, Albuquerque, NM, U.S.A.

08/1990 - 08/1993  — Assistant Professor, Department of Mathematics

& Statistics, University of New Mexico, Albuquerque, NM, U.S.A.

07/1987 - 07/1990  — G. C. Evans Instructor, Rice University,

Department of Mathematics, Houston, TX, U.S.A.

09/1984 - 06/1987  — Research Assistant at the Yang’s Institute,

SUNY@Stony Brook, NY, U.S.A.

09/1982 - 08/1984  — Assistant faculty member at the Institute

for Theoretical Physics, University of Wrocław, Wrocław, Poland

• Concurrent Visiting Positions (one semster or longer)

01.01-12.31 • 2004  — Max-Planck-Institut für Mathematik

Bonn, Germany (planned future visit)

01.01-12.31 • 1997  — Max-Planck-Institut für Mathematik

Bonn, Germany

01.01-05.30 • 1991  — School of Mathematics, Institute for Advanced Study,

Princeton, NJ, U.S.A.

• Other Visiting Positions (2 weeks to 2 months)

2006, 2005, 2003, 2001,  — Max-Planck-Institut für Mathematik
1999, 1995, 1993, 1989  Bonn, Germany
2005  — Institut de Mathématiques Élie Cartan,

Université Henri Poincaré Nancy 1, Nancy, France

2005, 2003, 2001, 1998  — Univeristà di Roma “La Sapienza”, Rome, Italy
2005, 2002, 2000  — Institut des Hautes Études Scientifiques,

Bures-sur-Yvette, France

2005, 2004, 2002, 1999, 1994 — Erwin Schrödinger Institute for Mathematical

Physics, Vienna, Austria

2005, 2001, 1995  — Humboldt Universität, Berlin, Germany

1997, 1996  — International Center for Mathematical Sciences,

Ediburgh, Scotland, United Kingdom.

1997  — Vrije Universitet, Amsterdam, Holland
1989  — Odense University, Odense, Denmark

• NSF Research Grants

2005 – 2008  — NSF Award in Geometric Analysis,

Einstein Metrics, Sasakian Geometry and Kähler Orbifolds; (with C. P. Boyer)

2002 – 2005  — NSF Award in Geometric Analysis,

Contact Geometry, Fano Orbifolds, and Einstein Metrics; (with C. P. Boyer)

1999 – 2002  — NSF Award in Geometric Analysis,

Contact Geometry and Einstein Metrics (with C. P. Boyer)

1995 – 1998  — NSF Award in Geometric Analysis,

Quaternionic Quotients, Einstein Manifolds, and the Topology of Moduli Spaces; (with C. P. Boyer and B. Mann)

1993 – 1995  — NSF Award in Geometric Analysis,

Quaternionic Quotients, Einstein Manifolds, and Mathematical Physics

1991 – 1993  — NSF Award in Geometric Analysis,

Geometry of Quaternionic Manifolds and Mathematical Physics

1989 – 1991  — NSF Award in Geometric Analysis,

Complex Geometry and Mathematical Physics; (with Y. S. Poon)

1988 – 1989  — NSF Award in Geometric Analysis,

Geometry and Mathematical Physics; (with Y. S. Poon and R. Wells)

• Other Grants and Awards

2005, 2003, 2001, 1998  — Istituto Nazionale di Alta Matematica (INdAM)

grants to support my visits at “La Sapienza”

2003  — National Research Council COBASE Grant

2000  — UNM A&S Senior Professor Research Semester Award
1995 – 1996  — Efroymson Award
1994 – 1997  — UNM Regent’s Lecturer Award

• Invited Addresses and Conferences

1. January 26, 2006 – Geometric Structures on 5-Manifolds, Ohio State University colloquium, Columbus, Ohio, USA.

2. December 16, 2005 – Geometric Structures on 5-Manifolds, Universität Köln geometry seminar, Köln, Germany.

3. November 22, 2005 – Geometric Structures on 5-Manifolds, Humboldt Universität geometry seminar, Berlin, Germany.

4. November 14, 2005 – Geometric Structures on 5-Manifolds, Jagiellonian University geometry seminar, Kraków, Poland.

5. November 1, 2005 – Seifert Bundles and Contact Structures, invited lecture at the Erwin Schrödinger Institute, Vienna, Austria.

6. October 10, 2005 – Manifolds with Positive Ricci Curvature, Arthur Besse seminar, École Polytechnique, Palaiseau, France.

7. September 23 – Sasakian Structures in Mathematics and Physics, invited lecture at the Géométrie Pseudo-Riemannienne et Supersymétrie, September 22-24, 2005; Institut de Recherche Mathématique Avancée de Strasbourg, Strasbourg, France.

8. September 1, 2004 – Geometric Structures on 5-manifolds, MPIM Oberseminar, MPIM, Bonn, Germany.

9. August 5 – Geometric Structures on 5-Manifolds, invited lecture at the SFB 478 workshop Curvature and Global Shape, 31 July – 7 August, 2005, Münster, Germany.

10. July 25, 2005 – Geometric Structures on 5-Manifolds, geometry seminar at the Università di Cagliari, Cagliari, Italy.

11. July 12, 2005 – Geometric Structures on 5-Manifolds, geometry seminar at the Università di Roma, “La Sapienza”, Rome, Italy.

12. July 3, 2005 – Geometric Structures on 5-manifolds, colloquium talk at Università di Bari, Bari, Italy.

13. June 7, 14, 2005 – Sasakian Geometry I-III, invited series of three lectures at the Institut Élie Cartan, Université Henri Poincaré, Nancy, France.

14. June 13, 2005 – Sasakian Geometry of 5-Manifolds, Arthur Besse seminar, École Polytechnique, Palaiseau, France.

15. December 10, 2004 – Einstein Metriken auf Exotischen Sphären (in German), Universität Hamburg, Hamburg, Germany.

16. November 19, 2004 – Einstein Metrics on Spheres, lecture given at the geometry seminar at the University of Bochum, Bochum, Germany.

17. November 2, 2004 – Einstein Metrics on Brieskorn Manifolds, lecture given at the geometry seminar at the University of Hamburg, Hamburg, Germany.

18. October 21, 2004 – Einstein Metrics on Spheres, lecture given at the geometry seminar at the University of Bonn, Bonn, Germany.

19. October 17, 2004 – Seifert Bundles, log del Pezzo Surfaces, and Einstein 5-Manifolds, MPI Oberseminar Topologie, Bonn, Germany.

20. October 12,14,15, 2004 – Sasakian Structures and Einstein Metrics I-III, series of three lectures at Centro di Ricerca Matematica Ennio De Giorgi, Differential Geometry and Topology, September 6 - November 12, 2004, Pisa, Italy.

21. May 12, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, invited colloquium, University of Bucharest, Bucharest, Romania.

22. April 22, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, invited colloquium, University of Göttingen, Göttingen, Germany.

23. April 6, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, invited lecture at the Erwin Schrödinger Institute in the program Geometric and analytic problems related to Cartan connections, Vienna, Austria.

24. March 18, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, MPIM Oberseminar, MPIM, Bonn, Germany.

25. February 28, 2004 – Einstein Metrics on Spheres, invited lecture at the Positive scalar curvature and related topics conference at LMU München, München, Germany.

26. February 2, 2004 – Transverse Fano Structures and Einstein Metrics on Spheres, geometry seminar at the University of Oxford, Oxford, England.

27. January 18-24, 2004 – Contact Structures, Complex Singularities, and Einstein Metrics I-III, a series of three invited plenary lectures at the XXIV Winter School in Geometry and Physics, Srni, Chech Republic.

28. July 21, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, geometry seminar at the Università di Cagliari, Cagliari, Italy.

29. July 4, 2003 – Negative Self-Dual Einstein Metrics as Quotients, lecture given at Università di Roma, “La Sapienza”, Roma, Italy.

30. June 24, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, geometry seminar at the Università di L’Aquila, L’Aquila, Italy.

31. May 2, 2003 – Transverse Fano Structures, and Einstein Metrics, geometry seminar at the University of California, Riverside, California, USA; May 2, 2003.

32. April 30, 2003 – Transverse Fano Structures and Positive Ricci Curvature, colloquium at the University of California, Riverside, California, USA; April 30, 2003.

33. April 8, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, colloquium at the Oregon State University, Corvalis, Oregon, USA.

34. April 8, 2003 – Transverse Fano Structures, Ricci Curvature, and Einstein Metrics, geometry seminar at the University of Oregon, Eugene, Oregon, USA.

35. November 22, 2002 – Transverse Fano Structures, Positive Ricci Curvature, and Einstein Metrics, invited lecture at the Erwin Schrödinger Institute, Vienna, Austria.

36. June 15, 2002 – On the Geometry of Sasakian-Einstein 5-Manifolds, invited lecture at the Joint International AMS/UMI Meeting in Pisa, June 12-16, Pisa, Italy.

37. June 6, 2002 – Sasakian Manifolds and Einstein Metrics in Dimension 5, invited lecture at the EDGE meeting on Special and Low-dimensional Geometries, June 3-7, Luminy, France.

38. May 28, 2002 – Contact Structures, Einstein Metrics and Positive Ricci Curvature, invited lecture at Institut Élie Cartan, Université Henri Poincaré, Nancy, France.

39. May 27, 2002 – Sasakian Manifolds and Einstein Metrics, Arthur Besse seminar, École Polytechnique, Palaiseau, France.

40. December 6, 2001 – Metric Cones, Holonomy, and Special Metrics, the SFB 288 Differentialgeometrie und Quantenphysik Kolloquium, Freie Universität Berlin, Berlin, Germany.

41. November 24, 2001 – Contact Structures and Algebraic Geometry, invited lecture at the conference Recent trends in mathematics, November 23-25, 2001, Schloss Klink, Germany.

42. November 19, 2001 – On Sasakian-Einstein 5-Manifolds, Humboldt Universität geometry seminar, Berlin, Germany.

43. September 26, 2001 – New Einstein Metrics in Dimensions Five and Seven, lecture given at the Università di Roma, “La Sapienza”, Rome, Italy.

44. August 4, 2001 – 3-Sasakian Geometry, Nilpotent Orbits, and Exceptional Quotients, LMS Durham Symposium, Special Structures in Differential Geometry, University of Durham, July 30 – August 9, 2001, Durham, England.

45. June 11, 2001 – Algebraic Geometry and Einstein manifolds in Dimension Five, invited lecture at Mathematisches Forschunginstitut Oberwolfach, Differentialgeometrie im Grossen, June 11 – 15, Oberwolfach, Germany.

46. May 22, 2001 – Algebraic Geometry and Compact Einstein 5-Manifolds, lecture given at the geometry seminar at the University of Bonn, Bonn, Germany.

47. March 21, 2001 – Sasakian-Einstein Geometry of 5-Manifolds, lecture given at the Università di Roma, “La Sapienza”, Rome, Italy.

48. May 19, 2000 – New Einstein Metrics on 5-Manifolds, Arthur Besse seminar, Ecole Polytechnique, Palaiseau, France.

49. August 29, 1999 – Killing Spinors, Einstein Manifolds, and Contact Geometry, invited lecture at the conference on Generalized Dirac Operators and Their Geometric Scattering Theory, Warsaw, 23 August - 4 September, 1999; International Stefan Banach Center of the Polish Academy of Sciences, Warsaw, Poland.

50. July 13, 1999 – Sasakian-Einstein Manifolds, Killing Spinors, and Near Horizon Geometries, Humboldt Universität geometry seminar, Berlin, Germany.

51. June 17, 1999 – Contact Structures, Einstein Manifolds, and Near Horizon Geometries, Universität München geometry seminar, München, Germany.

52. February 22, 1999 – Cones, Holonomy, and Einstein Manifolds, colloquium talk at West Washington University, Bellingham, Washington, USA.

53. June 18, 1998 – New Families of Einstein 7-manifolds, colloquium talk at Università di Bari, Bari, Italy.

54. June 2,4,9,2, 1998 – A Short Introduction to 3-Sasakian Manifolds I-IV, four lectures given at the Università di Roma, “La Sapienza”, Rome, Italy.

55. December 4, 1997 – Quaternionic Geometry and Einstein Manifolds, colloquium given in the University of Leipzig, Leipzig, Germany.

56. November 9, 1997 – Hypercomplex Structures on Toral Quotients, talk given in the Special Session of the AMS Meeting, Albuquerque, New Mexico, USA.

57. October 9, 1997 – Einstein Manifolds From Spheres, report on my research presented to the Fachbeirat at the Max-Planck-Institut für Mathematik in Bonn, Bonn, Germany.

58. August 18, 1997 – Quaternionic Geometry and Einstein Manifolds, lecture at the ICMS conference Complex Methods in Differential Geometry 18-29 August, 1997, Edinburgh, Scotland.

59. July 15, 1997 – Quaternionic Geometry and Einstein Manifolds, lecture given in the MPI Faltings-Harder-Manin-Zagier Oberseminar, MPI, Bonn, Germany.

60. June 9, 1997 – Making Einstein Manifolds Out of Spheres, geometry seminar at the Jagiellonian University, Kraków, Poland.

61. April 7, 1997 – On Betti Numbers of 3-Sasakian Manifolds, Topologie Oberseminar MPI, Bonn, Germany.

62. February 21, 1997 – Compact Einstein 7-manifolds of Positive Scalar Curvature and Arbitrary Second Betti Number, and 3-Sasakian Geometry and Killing Spinors, two lectures at the Meeting on Dirac Operators and Geometry, Warsaw, 16 February – 1 March, 1997, International Stefan Banach Center of the Polish Academy of Sciences, Warsaw, Poland.

63. February 4, 1997 – Compact Einstein 7-manifolds of Positive Scalar Curvature and Arbitrary Second Betti Number, geometry seminar at the University of Bonn, Bonn, Germany.

64. November 13, 1996 – Making Einstein Manifolds Out of Spheres, colloquium at the University of California at Riverside, Riverside, California, USA.

65. June 4,6,27, 1996 – Geometry and Topology of 3-Sasakian Manifolds I-III, three lectures at the ICMS Workshop Interplays Between Geometry and Topology, June 3-29, 1996, Edinburgh, Scotland.

66. April 14, 1996 – Hypercomplex manifolds and 3-Sasakian geometry, contributed paper in a Special Session at the AMS meeting, April 13-14, 1996, New York, New York, USA.

67. December 1, 1995 – On Betti Numbers of 3-Sasakian Manifolds, contributed paper in a Special Session at the second joint meeting of AMS and SMM, November 29 – December 2, 1995, Guanajuato, Mexico.

68. November 21, 1995 – Geometry and Topology of 3-Sasakian Manifolds: On Betti Numbers of 3-Sasakian Spaces, Humboldt Universität geometry seminar, Berlin, Germany.

69. November 16, 1995 – Quaternionic Geometry, Supergravity, and Einstein Manifolds, colloquium at Augsburg University, Augsburg, Germany.

70. November 14, 1995 – Geometry and Topology of 3-Sasakian Manifolds: Quotient Construction, Humboldt Universität geometry seminar, Berlin, Germany.

71. November 9, 1995 – Quaternionic Geometry, 3-Sasakian Manifolds, and Supergravity, the SFB 288 Differentialgeometrie und Quantenphysik Kolloquium, Freie Universität Berlin, Berlin, Germany.

72. June 12, 1995 – 3-Sasakian Manifolds, invited lecture at Mathematisches Forschunginstitut Oberwolfach, Differentialgeometrie im Grossen, June 12 – 16, Oberwolfach, Germany.

73. June 9, 1995 – Geometry and Topology of 3-Sasakian Manifolds, geometry seminar at the Max-Planck-Institut, Bonn, Germany.

74. May 22, 1995 – Geometry and Topology of 3-Sasakian Manifolds, invited lecture in Arthur Besse Seminar at École Polytechnique, Paris, France.

75. January 10, 1995 – 3-Sasakian Manifolds, lecture presented at the conference on Representation Theory and Differential Geometry, University of Adelaide, Adelaide, Australia.

76. September 28, 1994 – Einstein Metrics and 3-Sasakian Geometry, lecture presented at Erwin Schrödinger Institute, Vienna, Austria.

77. September 9, 1994 – On 3-Sasakian Manifolds, lecture presented at the Meeting on Quaternionic Structures in Mathematical Physics, September 5 – 9, 1994; SISSA, Trieste, Italy.

78. November 17, 1993 – On Strongly Inhomogeneous Einstein Manifolds with Positive Scalar Curvature and 3-Sasakian Structure, talk presented at Columbia University, New York, New York, USA.

79. November 16, 1993 – On Strongly Inhomogeneous Einstein Manifolds with Positive Scalar Curvature and 3-Sasakian Structure, talk presented at SUNY at Stony, Brook, Stony Brook, New York, USA.

80. November 15, 1993 – On Strongly Inhomogeneous Einstein Manifolds with Positive Scalar curvature and 3-Sasakian structure, talk presented at the Institute for Advanced Study, Princeton, New Jersey, USA.

81. July 1, 1993 – A Construction of Strongly Inhomogeneous Einstein Manifolds with 3-Sasakian Structure, talk presented at the Max-Planck Institut Harder-Hirzebruch-Zagier Oberseminar, Max-Planck Institut für Mathematik, Bonn, Germany.

82. June 4, 1993 – Einstein Manifolds in Field Theories with Supersymmetry, talk given at the University of Wrocław, Wrocław, Poland.

83. May 26, 1992 – O Pewnych Siedmiowymiarowych Rozmaitościach Einsteina, Wrocław University, Wrocław, Poland.

84. February 29, 1992 – Quaternionic Reduction and Einstein Manifolds, The New Mexico Geometry and Topology Symposium, Las Cruces, New Mexico, USA.

85. February 12, 1991 – Group Actions, Quotients, and Quaternionic Manifolds, invited talk at SUNY@Stony Brook, Stony Brook, New York, USA.

86. January 27, 1991 – Group Actions, Quotients, and Quaternionic Manifolds, talk given at The Institute for Advanced Study, Princeton, New Jersey, USA.

87. May 14, 1990 – Self-dual Einstein Metrics with Nonzero Scalar Curvature, talk given at The University of Warwick, England.

88. August 25, 1989 – Quotient Constructions of Special Riemannian Geometries, colloquium at Odense University, Odense, Denmark.

89. August 23, 1989 – Quotient Constructions of Special Riemannian Geometries, seminar talk at the Danmarks Tekniske Høskole, Copenhagen, Denmark.

90. June 16, 1989 – Generalized Symplectic Reduction, seminar talk at the Max-Planck-Institut, Bonn, Germany.

91. February 21, 1989 – Generalized Symplectic Reduction, invited colloquium at University of California, Santa Barbara, California, USA.

92. December 8, 1988 – New Metrics with Sp(n) · Sp(1) Holonomy, invited colloquium at Tulane University, New Orleans, Louisiana, USA.

PUBLICATIONS

• Articles published or accepted for publication

1. On η-Einstein Sasakian Geometry, Communications in Mathematical Physics 262 (2006), 177-208; (with C. P. Boyer and P. Matzeu).

2. Sasakian geometry Einstein metrics on spheres, Short Program on Riemannian Geometry, 47-61, CRM Proceedings & Lecture Notes, 40, AMS, Providence, RI, 2005; (with C. P. Boyer).

3. Einstein metrics on spheres, Annals of Mathematics 162 (2005), 557-580; (with C. P. Boyer and J. Kollár).

4. Einstein metrics on exotic spheres in dimension 7,11, and 15, Experimental Mathematics 14 (2005), 61-66; (with C. P. Boyer, J. Kollár, and E. Thomas).

5. Sasakian Geometry, Hypersurface Singularities, and Einstein Metrics, Rendiconti del Circolo Matematico di Palermo (2), Suppl. No. 75 (2005), 57-87; (with C. P. Boyer).

6. Toric self-dual Einstein metrics as quotients, Communications in Mathematical Physics 253 (2005), 337-370; (with C. P. Boyer, D. M. J. Calderbank, and P. Piccinni).

7. On Positive Sasakian Geometry, Geometriae Dedicata 101 (2003), 93-102; (with C. P. Boyer and M. Nakamaye).

8. New Einstein Metrics on 8#(S^{2} × S^{3}), Differential Geometry and Its Applications 19 (2003); 245-251; (with C. P. Boyer).

9. The Geometry of Sasakian-Einstein 5-Manifolds, Mathematische Annalen 325 (2003), 485-524; (with C. P. Boyer and M. Nakamaye).

10. Sasakian Geometry, Homotopy Spheres, and Positive Ricci Curvature, Topology 42 (2003), 981-1002; (with C. P. Boyer and M. Nakamaye).

11. Einstein Metrics on Rational Homology 7-Spheres, Annales de l’Institut Fourier 52 (2002), 1569-1584; (with C. P. Boyer and M. Nakamaye).

12. Rational Homology 5-Spheres with Positive Ricci Curvature, Mathematical Research Letters 9 (2002), 521-528; (with C. P. Boyer).

13. Sasakian-Einstein Structures on #9(S^{2} × S^{3}), Transactions of the American Mathematical Society 354 (2002), 2983-2996; (with C. P. Boyer and M. Nakamaye).

14. 3-Sasakian Geometry, Nilpotent Orbits, and Exceptional Quotients, Annals of Global Analysis and Geometry 21 (2002), 85-110; (with C. P. Boyer and P. Piccinni).

15. New Einstein Metrics in Dimension Five, Journal of Differential Geometry 57 (2001); 443-463; (with C. P. Boyer).

16. Contact Geometry and Einstein Manifolds, Proceedings of the American Mathematical Society 129 (2001), 2419-2430; (with C. P. Boyer).

17. On Sasakian-Einstein Geometry, International Journal of Mathematics 11 (2000), 873-909; (with C. P. Boyer).

18. A Note on Contact Toric Geometry, Journal of Geometry and Physics 35 (2000), 288-298; (with C. P. Boyer).

19. 3-Sasakian Manifolds, Encyclopaedia of Mathematics, Supplement II, pp. 449-451, Kluwer Academic Publishers, (2000).

20. 3-Sasakian Manifolds; Surveys in Differential Geometry VI: Essays on Einstein Manifolds A supplement to the Journal of Differential Geometry, pp. 123-184, (eds. C. LeBrun, M. Wang); International Press, Cambridge (1999); (with C. P. Boyer).

21. Quaternionic geometry and 3-Sasakian manifolds, Quaternionic structures in mathematics and physics (Trieste, 1994), 7–24 (electronic), Int. Sch. Adv. Stud. (SISSA), Trieste, (1998); (with C. P. Boyer and B. Mann).

22. Hypercomplex Structures From 3-Sasakian Structures, Journal für die reine und angewandte Mathematik 501 (1998), 115-141; (with C. P. Boyer and B. Mann).

23. A Note on Smooth Toral Reductions of Spheres, Manuscripta Mathematica 95 (1998), 149-158. (with C. P. Boyer and B. Mann).

24. Compact 3-Sasakian 7-Manifolds with Arbitrary Second Betti Number, Inventiones Mathematicae 131 (1998), 321-344; (with C. P. Boyer, B. Mann, and E. Rees).

25. The Twistor Space of a 3-Sasakian Manifolds, International Journal of Mathematics 8 (1997), 31-60; (with C. P. Boyer).

26. Einstein Manifolds of Positive Scalar Curvature with Arbitrary Second Betti Number, Balkan Journal of Geometry and Its Applications 1, no. 2 (1996), 1-8. (with C. P. Boyer, B. Mann, and E. Rees).

27. On Betti Numbers of 3-Sasakian Manifolds, Geometriae Dedicata 63 (1996), 45-68; (with S. Salamon).

28. On Strongly Inhomogeneous Einstein Manifolds, Bulletin of the London Mathematical Society 28 (1996), 401-408; (with C. P. Boyer and B. Mann).

29. Hypercomplex Structures on Stiefel Manifolds, Annals of Global Analysis and Geometry 14 (1996), 81-105; (with C. P. Boyer and B. Mann).

30. Some New Examples of Inhomogeneous Hypercomplex Manifolds, Mathematical Research Letters 1 (1994), 531-538; (with C. P. Boyer and B. Mann).

31. Geometry and topology of 3-Sasakian manifolds, Journal für die reine und angewandte Mathematik 455 (1994), 183-220; (with C. P. Boyer and B. Mann).

32. New Examples of Inhomogeneous Einstein Manifolds of Positive Scalar Curvature, Mathematical Research Letters 1 (1994), 115-121; (with C. P. Boyer and B. Mann).

33. 3-Sasakian Manifolds, Proceedings of Japan Academy vol. 69, Ser. A, (1993), 335-340; (with C. P. Boyer and B. Mann).

34. Quaternionic Quotients and Einstein Manifolds, in Communications in Analysis and Geometry vol. 1, no. 2, (1993), 1-51; (with C. P. Boyer and B. Mann).

35. Nonzero Scalar Curvature Generalizations of the ALE Instantons, Journal of Mathematical Physics 33 (1992), 1765-1771; (with T. Nitta).

36. Geometry of the Scalar Couplings in N = 2 Supergravity Models, Classical and Quantum Gravity 9 (1992), 27-40.

37. Multicenter Metrics with Negative Cosmological Constant, Classical and Quantum Gravity 8 (1991), 1529-1543.

38. Duality and Yang-Mills Fields on quaternionic Kähler manifolds, Journal of Mathematical Physics 32 (1991), 1263-1268; (with Y. S. Poon).

39. Quaternionic Reduction and Quaternionic Orbifolds, Mathematische Annalen 282 (1988), 1-21; (with B.H. Lawson, Jr.)

40. A Generalization of Momentum Mapping Construction for Quaternionic Kähler Manifolds, Communications in Mathematical Physics 108 (1987), 117-138.

41. New Matter Couplings in N=2 Supergravity, Nuclear Physics B289 (1987), 573-588.

42. Hyperkähler and Quaternionic Kähler Nonlinear Sigma Models, Nuclear Physics B271 (1986), 402-416.

43. CP (n) nonlinear sigma models with holomorphic constraints in d=2,4 dimensions, Physical Reviews D29 (1984), 1809-1818.

• Articles submitted for publication

1. Einstein metric on rational homology spheres, submitted to the Journal of Differential Geometry; (with C. P. Boyer).

2. Highly connected manifolds with positive Ricci curvature, submitted to Geometry and Topology; (with C. P. Boyer).

• To be submitted and in preparation

1. Constructions in Sasakian Geometry, in preparation; (with C. P. Boyer and L. Ornea).

2. Sasakian Manifolds, research monograph for the Oxford University Press; in preparation (with C. P. Boyer).

3. On Sasakian Futaki Invariants, in preparation; (with C. P. Boyer and S. Simanca).

4. Sasakian Geometry in Five Dimensions, in preparation; (with C. P. Boyer).

Last updated: February 1, 2006